OpenStudy (4meisu):

Use the various laws of logarithims to find the values of the positive integers a and b: 3log2(a )+ bLog2(4)-log2(64)=1

OpenStudy (anonymous):

can I help

OpenStudy (4meisu):

yes please!

OpenStudy (anonymous):

what do you think it is

OpenStudy (4meisu):

5(a)(i) for reference :)

OpenStudy (4meisu):

Well so far I've simplified it to log2(a^3+4^b/64)=1, but I'm not sure what to do next

OpenStudy (4meisu):

I've also got 2^(64-a^3) = 4^b in my working but idk where to go from there

OpenStudy (anonymous):

great what is this from a book or flvs

OpenStudy (anonymous):

@4meisu

OpenStudy (anonymous):

are you there

hartnn (hartnn):

` log2(a^3+4^b/64)=1` how did you get a `+` sign in between?

OpenStudy (4meisu):

oh yep, you're right hartnn

OpenStudy (priyar):

do u know that u can multiply the value when bases are same ?

OpenStudy (priyar):

@4meisu ?

OpenStudy (priyar):

and..here u have base as 2 on the LHS.. so try to write 1 as log something.base 2 OK?

OpenStudy (priyar):

@4meisu ? are u there?

hartnn (hartnn):

so you got till here? \(\log_2(\dfrac{a^34^b}{64})=1 \) ?

OpenStudy (4meisu):

Yep, I got till there

OpenStudy (priyar):

now we can write 1 as \[\log_{2} 2\]

OpenStudy (priyar):

do u know that?

hartnn (hartnn):